• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Xiaofeng

    Log entry time 08:01:47 on November29,2001

    Entry number 78264

    keyword=Shift summary

    Shift of Thurs 11/29, 00h -> 08h

    00h05 : shift begins, no beam.
    00h15 : MCC calls to say they will restore current begining with a low value and then increasing to 55 microA
    00h25 : run 2635 starts. junk, wrong prescale factor
    00h31: run 2636 starts with new prescale factor: ps1=5200,ps2=212,ps3=2460,ps4=268,ps5=1,ps8=65535
    02h20: MCC need to check the main frame, beam stopped
    02h33: Beam is back
    03h01: MCC RF tripped, and phoned about QE measurement for 15~20 minutes
    03h15 : beam is back
    03h30: MCC called about the beam trip in HallC's orbit, so they sent turned beam
    03h35 : left arm Q1 tripped, asked for controlled access, and reset it
    04h30: asked for beam permit
    05h00: left arm Q1 tripped again, asked for controlled access and reset it
    06h20: asked MCC to switch HallA to beam permit. However, before sending beam to us , MCC needed at leat 30 minutes to fix their problems.
    06h30: left arm Q1 tripped again.
    07:05am MCC fixed their problem. We asked for controller access , and wait for expert to fix the problem.
    07:55am Ed Flots came to help.

    run: events: Q: sumQ:

    2635 junk
    2636 2.7 22.9 1988.4
    2637 2.7 23 2011.4
    2638 2.7 23 2034.4
    2639 1.0 8.6 2043
    2640 1.48 9.75 2052.7
    2641 beam not stable